# Monitoring of water volume in a porous reservoir using seismic data: Validation of a numerical model with a field experiment

### v3

@@ 4 Neural network-based characterization of water storage
4.1 Noise model and source function normalization
@@
where $\bm{m}$ contains all the physical and geometrical parameters of the model and $e$ accommodates additive noise components. The forward operator $\mathcal{A}$ is used to map the model parameters $\bm{m}$ to the synthetic seismic data vector $X$, simulated using the coupled viscoelastic-poroviscoelastic material model by the DG method in three spatial dimensions [5, 14].
To simulate the noise of the field measurements, we apply the following noise model
@@
where $\epsilon^{A}$ and $\epsilon^{B}$ represent independent zero-mean Gaussian random variables. $|X|_{\max}$ is the maximum absolute value of $X$. The two noise components correspond to additive white noise and amplitude-related noise, respectively, contributing to a diverse range of noise levels. By adjusting $A$ and $B$ within the intervals $[0.05,2]$ per cent and $[0,2]$ per cent, respectively, wide noise variations are introduced.
As discussed in Section 2.2, the field data was measured for 3 seconds. We estimated the standard deviation of the Gaussian white noise component by analyzing the measured data within a time window spanning from 1 second to 2 seconds. Subsequently, we compared this estimated value to the maximum amplitudes observed in the traces. This gave an approximation of the noise level in the real data ($A_{\rm{meas}}\approx 0.4$) per cent and the applied noise levels in model (5) covers such value. The selection for parameter $B$ interval is arbitrary.
In order to obtain a source-independent inversion, we use a deconvolution operation to remove the effect of the seismic source time function [16]. We transform the transient signals to the frequency domain and use data from a reference point, which is an additional receiver (see Fig. 2), as the system response function. This allows us to change the original observation model (Eq. (4)) with a new formulation
$\frac{F_{\Upsilon}(\omega_{\ell})}{F_{\Upsilon_{\rm ref}}(\omega_{\ell})}$ $=$ $\frac{F_{X}(\omega_{\ell})}{F_{X_{{\rm ref}}}(\omega_{\ell})}+\hat{e}_{\ell},\quad\ell=1,\ldots,N_{f}$ (6)
@@
where $F_{\{\Upsilon,\Upsilon_{\rm ref},X,X_{{\rm ref}}\}}$ represents the data in the frequency domain obtained through the Fourier transform. The subscript ${\rm ref}$ denotes the data corresponding to the reference point. $\hat{e}$ is the noise in the frequency domain formulation after normalization, $\omega_{\ell}$ is the frequency at the $\ell$th index, and $N_{f}$ indicates the total number of frequencies. Note that the model (7) is applied separately to both velocity components, $v_{\rm s}$ and $w_{\rm s}$, at each receiver. Since (7) may become ill-conditioned when the denominator approaches zero, we adopt a Wiener filtering approach for regularization, following the method applied in [23].
4.2 Examples of real and synthetic seismic data
Figure 3 presents a sample shot gather from the real dataset, displaying the horizontal $v_{\rm s}$ and vertical $w_{\rm s}$ velocity components along with the reference records as an example in the top right corner. The bottom row shows the real and imaginary parts of the Fourier transformed and source function normalized records, which serve as inputs to the neural network. For the shown data, water table level was at $-88.7$ cm.
Figure 3: An example of the measured data. The figures on the left show the horizontal $v_{\rm s}$ component, and the figures on the right the vertical $w_{\rm s}$ component. The velocity components on the reference point are shown to illustrate the data behaviour.
Figure 4 shows an example of the simulated data. Here the water table level was at $-72.10$ cm and the noise level parameters were set to $A=1$ per cent and $B=1$ per cent. The discrepancies between the measured and synthetic seismograms in these figures arise from the random selection of material parameters used to generate the synthetic data, as well as differences in the water table levels between the real and synthetic datasets.
Figure 4: An example of the synthetic data. The pictures on the left show the horizontal $v_{\rm s}$ component, and the figures on the right the vertical $w_{\rm s}$ component. The reference data traces on the top panel are shown for the noisy and noiseless data for both velocity components.
@@ 4.3 Neural networks
We generated 15,000 training and 3,000 validation samples using fifth-order polynomials and the source wavelet (1) with the DG solver. The values for the model input $\bm{m}$ were independently drawn from the propability distributions described in Sections 3.1 and 3.2 and used as inputs to the forward model $\mathcal{A}$ (see Eq. (7)). The inviscid material model-based element size criterion is approximately 1.9 elements per shortest wavelength for the samples in the validation dataset and 2.0 for training dataset samples.
A fully connected neural network was employed in this study to determine the water volume using seismic data. By training on data with randomized model parameters $\bm{m}$ (see Eq. (7)), the network learns to focus on the relevant seismic signatures of water volume while marginalizing the influence of less significant factors. We down-sample the synthetic data to a sampling frequency of 4 kHz, then generate five copies of the clean training and validation datasets, each corrupted with Gaussian noise according to the noise model in (5). The noise corrupted time domain data are transformed to the frequency domain according to model (7) and denoted as $\hat{X}_{\textrm{noised}}$ in the following. The real and imaginary components of $\hat{X}_{\textrm{noised}}$ are used as inputs to the network. We select $N_{f}=21$ frequencies between 30 to 90 Hz, see Fig. 4 as an example.
The training process is implemented using the TensorFlow [2] and Keras [4] libraries. We employ the Adam optimizer [15] and utilize the RandomSearch algorithm from the Keras Tuner library [20] to optimize the model’s performance by exploring different hyperparameter configurations. The Keras Tuner is allowed to randomly select the activation function, learning rate, number of hidden layers, number of neurons per layer, and $L_{2}$ regularization penalty factor. The activation functions considered are “relu,” “sigmoid,” “tanh,” “selu,” “swish,” and “LeakyReLU,” while the number of hidden layers ranges from one to six. The number of neurons per layer varies between 100 and 5,000. The learning rate is selected from 1e-3, 1e-4, 1e-5, and the $L_{2}$ regularization penalty is chosen from 1e-5, 1e-6, 1e-7, 1e-8. For all networks in this study, the batch size is set to 256.
@@
The best-performing network is selected based on MSE for validation. After testing various hyperparameter combinations, the final network consists of five hidden layers with 2,570 (layer 1), 3,920 (layer 2), 3,360 (layer 3), 2,730 (layer 4), and 3,400 (layer 5) neurons. It employs the LeakyReLU activation function, a learning rate of 1e-5, and an $L_{2}$ regularization penalty factor of 1e-5. The output layer uses a linear activation function, and early stopping is enabled during both hyperparameter tuning and training to enhance generalization and prevent overfitting.
Figure 5 illustrates the approach for estimating water volume from seismic data. The figure shows the sequential steps in data processing, followed by the neural network model, which consists of five hidden layers. Additionally, it includes the learning curves for both training and validation, highlighting the model’s convergence and generalization during training.
Figure 5: Graphical illustration of the process for estimating water volume from seismic data. The left side illustrates the pre-processing phase, consisting deconvolution operation to normalize the source wavelet effects. On the right, the neural network-based estimation phase illustrates the applied network architecture. Each neuron in the hidden layers is represented by the symbol $a$, indicating its activation. In the architecture, $\{m_{1},m_{2},\ldots,m_{5}\}$ denote the number of neurons in each hidden layer. The loss curves as a function of epochs for the training and validation databases are shown in the top right corner.
@@ 5 Results
All results, i.e. the simulation of wave fields and estimation of water volumes via neural networks, were computed using the computer cluster Puhti at the CSC – IT Center for Science Ltd, Finland. A detailed description of the supercomputer Puhti can be found from the CSC’s website [1]. Computational grids used in this work were build using COMSOL Multiphysics.
@@ 5.1 Predictions of water volume
We test the applicability of the trained neural network model via two different test dataset. These datasets are defined as
@@
  • Field Measurements: Field data were acquired at seven distinct water table levels, ranging from a depth of –31.3 cm to –88.7 cm. At each level, three repeated measurements were recorded for each of three different drop heights, resulting in a dataset comprising $3\times 3\times 7=63$ observations. Using our knowledge of the actual water table level and the geometry of the sand pool, we can calculate the volume of the water-saturated zone. Additionally, based on a previous study [21], we assume the nominal porosity of the material to be 35 per cent.
@@
  • Synthetic Dataset: This database contains a total of 3,000 samples. We used the same prior to randomize material parameters as in the training and validation databases. Instead of randomizing the noise parameters $A$ and $B$ in (5), we set them to $1$ per cent each, representing moderate noise levels. To introduce varying numerical noise in the data compared to the training and validation databases, we used a mesh density criteria of 2.5 elements per wavelength, fourth-order basis functions, and a Ricker wavelet as the source function, defined as
    $g=(1+2b(t-t_{0})^{2})\exp\left(b\left(t-t_{0}\right)^{2}\right).$ (9)
    The parameter $b$ and time delay $t_{0}$ are defined as for wavelet (1).
@@
The Keras Tuner-optimized network architecture undergoes ten training runs, and the final water volume per sample result is determined as the average of these ten estimates. Figure 6 presents a comparison between the estimated (average) water volumes and their corresponding true values. The results for all drop heights are shown with different colors, red color denotes the repeated measurements from drop height 5 cm ($\rm{dh}_{5}^{1,2,3}$), green drop height 10 cm ($\rm{dh}_{10}^{1,2,3}$), and blue drop height 15 cm ($\rm{dh}_{15}^{1,2,3}$). To get a crude approximation for the uncertainty, we assumed that the porosity value in the field measurements database is uncertain in a sense that we assumed $\phi_{\rm{true}}\in[0.95,1.05]\cdot 35$ per cent, that can be used to compute the error bars shown for each estimate with real data. The figure shows also the estimates for the synthetic dataset. These results demonstrate the potential of using proposed neural network-based approach to recover the water volume.
@@
Figure 6: Left: Estimated volumes of water as a function of the true value. Red denotes repeated measurements from drop height 5 cm ($\rm{dh}_{5}^{1,2,3}$), green from drop height 10 cm ($\rm{dh}_{10}^{1,2,3}$), and blue from drop height 15 cm ($\rm{dh}_{15}^{1,2,3}$). Right: Histogram of relative prediction errors, with the color coding of the asterisks corresponding to the same drop heights as shown on the left.
@@
The results at water table level -36.2 cm reveal a clear outlier in Fig. 6. To analyze the differences between repeated measurements, we utilize the RMSE. Let $d_{i}^{k}$ represent the normalized seismic data (see model (7)) for the $k$’th repeated measurement at drop heights $i\in\{5,10,15\}$ in the frequency domain, with the real and imaginary components stacked. The root mean square error (RMSE) error can now be expressed as follows:
@@
Table 3: Root mean square errors (10) for different data combinations for measurements taken at a water table level of -36.2 cm.
@@ 5.2 SHAP analysis
6 Discussion and Conclusions
In this study, we investigated water volume estimation using both field and synthetic seismic data. The field measurements were done for a reservoir for which the water table level is controllable and known. Furthermore, we possessed accurate knowledge of the reservoir’s geometry and physical characteristics. Our approach involved testing the field data against a proposed neural network model, developed through a training phase utilizing synthetic data.
The physical model for synthetic wave propagation computations was a coupled poroviscoelastic–viscoelastic system. The wave propagation problem was solved on a GPU cluster using the nodal DG method coupled with the Adams-Bashforth time-stepping scheme to generate synthetic seismograms. In the synthetic model, the material was assumed to be homogeneous, and the water table level was allowed to move freely within the geometry.
For the inverse problem, we employed a fully connected neural network to estimate water volume from seismic data. The databases used during the training phase were synthetically generated using a wave propagation solver. The training data were corrupted with noise at levels similar to those found in real data. In our proposed approach, the source wavelet is unknown, and we employ a deconvolution-based method to normalize the source function. The input data for the neural network consisted of seismic data in the frequency domain.
A crucial element of our research was the examination of how closely the estimations from seismic data, based on neural networks, matched with the actual water volumes in the controlled reservoir environment. These correlations are key to verifying the accuracy of our neural network model. Quantitative comparison between the seismic data-derived estimates and the ground truth of the reservoir’s water volumes is fundamental to demonstrating the reliability of our approach. It is especially pertinent when considering the potential application of our methods in varied and uncontrolled real-world scenarios. It is expected that when encountering out-of-training-distribution samples, whether due to deviations in material parameters or geometric inaccuracies, the estimation accuracy may deteriorate. Additionally, certain parameters are likely to have a greater impact on estimation errors than others, making sensitivity to specific variables an important consideration in assessing the robustness of the model.
In this study, we computed the average absolute SHAP values at the receiver level for the field data. Since SHAP values provide insight into which of the receivers contributes the most to the estimates, we tested the estimation accuracy by selecting the ten most contributing receivers and repeating the estimation procedure. In this case, the network was re-optimized using Keras Tuner tools. Additionally, we tested the estimation accuracy by selecting receivers randomly. For the randomly selected receivers, we used the same network as was used for SHAP value-based selection. It was observed that the estimation accuracy was significantly affected when randomly selected receivers were used. On the other hand, the accuracy between the full array and SHAP value-based selection methods was similar. For synthetic data, the effect on estimation accuracy was not as dramatic when randomly selected receivers were used.
An important next step in research involves exploring larger test sites. From a methodological perspective, approaches to quantify and characterize the uncertainty in neural network-based estimates, as well as methods to normalize the source wavelet in the time domain, could be valuable directions. Furthermore, depending on the characteristics of the site under investigation, it may be necessary to extend the current physical models. These extensions could involve accommodating inhomogeneous and possibly anisotropic materials, addressing the challenges associated with regions of partial saturation, and incorporating more complex noise models. Additional work could also investigate which parameters have the greatest impact on uncertainty in model outputs. Moreover, expanding the synthetic databases to include multiple source locations, instead of relying on a single, arbitrarily selected source, could help address challenges associated with complex geometries and material inhomogeneities.

### v4

@@ 4 Neural network-based characterization of water storage
4.1 Noise model, source function normalization, and data examples
@@
where $\bm{m}$ contains all the physical and geometrical parameters of the applied model, such as porosity, tortuosity, and water-table level. The term $e$ represents additive noise components. The forward operator $\mathcal{A}$ is used to map the model parameters $\bm{m}$ to the synthetic seismic data vector $X$, simulated using the coupled viscoelastic-poroviscoelastic material model by the DG method in three spatial dimensions [6, 17].
To simulate measurement noise, we generate the noisy trace using the model
@@
where $\epsilon^{A}$ and $\epsilon^{B}$ represent independent zero-mean Gaussian random variables, and $|X|_{\max}$ is the maximum absolute value of $X$. The two noise components correspond to additive white noise and amplitude-related noise, respectively, contributing to a diverse range of noise levels. This decomposition is motivated by practical considerations: in datasets with large dynamic range, a pure white-noise model disproportionately affects weaker signals while leaving stronger signals largely unperturbed. By including an amplitude-related term, the noise level scales with the signal, producing a more realistic distribution across all traces. While the additive component can be estimated from the data, the amplitude-dependent component cannot be robustly inferred. Therefore, parameters $A$ and $B$ are varied over wide intervals ($A\in[0.05,2]\%$, $B\in[0,2]\%$) to explore different noise conditions.
The seismic measurements were recorded continuously, but each source excitation produced a 3-second-long data segment, which was downloaded from the recorders based on microsecond-accurate, GPS-timed source initiations. For the neural network input, a 0.35-second window was extracted from both field and simulated data. This duration ensures that all significant wave reflections and late arrivals have attenuated to negligible amplitudes, preventing artificial truncation of energy and allowing each trace to end near zero amplitude.
After the waves produced by the seismic source have attenuated, we estimated the standard deviation of the Gaussian white noise component by analyzing a time window between 1 and 2 seconds. Comparing the estimated noise amplitude to peak signal amplitudes yielded an approximate noise level of $A_{\rm meas}\approx 0.4\%$.
In order to obtain a source-independent inversion, we use a deconvolution operation to remove the effect of the seismic source time function [20]. We transform the transient signals to the frequency domain and use data from a reference receiver (see Fig. 2) as the system response function. This allows us to rewrite the observation model (Eq. 14) as
$\frac{F_{\Upsilon}(\omega_{\ell})}{F_{\Upsilon_{\rm ref}}(\omega_{\ell})}$ $=$ $\frac{F_{X}(\omega_{\ell})}{F_{X_{{\rm ref}}}(\omega_{\ell})}+\hat{e}_{\ell},\qquad\ell=1,\ldots,N_{f},$ (16)
@@
where $F_{\Upsilon}$ is the Fourier transform of the measured seismic data at a given receiver, $F_{\Upsilon_{\rm ref}}$ is the Fourier transform of the data at the reference receiver, $F_{X}$ denotes the Fourier transform of the simulated (noise-free) model response, and $F_{X_{\rm ref}}$ is the corresponding simulated response at the reference receiver. Here, $\hat{e}$ is the noise term in the frequency-domain formulation after normalization, $\omega_{\ell}$ is the frequency at index $\ell$, and $N_{f}$ is the total number of frequencies. Note that the model (17) is applied separately to both velocity components, $v_{\rm s}$ and $w_{\rm s}$, at each receiver.
In practice, the division in Eq. (17) is implemented using the conjugate of $F_{X_{\rm ref}}(\omega_{\ell})$, leading to a denominator of the form $|F_{X_{\rm ref}}(\omega_{\ell})|^{2}$. Since this denominator may approach zero, we apply a Wiener-type regularization. Following [31], the denominator $|F_{X_{\rm ref}}(\omega_{\ell})|^{2}$ is replaced by $|F_{X_{\rm ref}}(\omega_{\ell})|^{2}+\epsilon\langle|F_{X_{\rm ref}}(\omega_{\ell})|^{2}\rangle,$ where $\epsilon=0.1$ is a stabilization constant. This regularization is applied directly within the frequency-domain ratio in Eq. (17). It does not replace Eq. (17) nor introduce any additional filtering procedure, it only serves to avoid numerical instabilities when the reference spectrum has small amplitudes.
To illustrate the behavior of real seismic data, Fig. 4 shows a sample shot gather from the field dataset, displaying the horizontal $v_{\rm s}$ and vertical $w_{\rm s}$ velocity components along with reference traces. The bottom row presents the real and imaginary parts of the Fourier-transformed and source-normalized records, which serve as inputs to the neural network. For this example, the water table was at 88.7-88.7 cm.
Figure 4: Measured seismic data: horizontal $v_{\rm s}$ component (left), vertical $w_{\rm s}$ component (right). Reference traces are shown in the top right panels.
Similarly, Fig. 5 presents an example of synthetic data with a water table at 72.10-72.10 cm and noise parameters $A=1\%$, $B=1\%$. Differences from real data arise from randomly selected material parameters and differing water-table levels. These examples illustrate the correspondence between measured and simulated records, and the type of inputs provided to the neural network.
Figure 5: Synthetic seismic data: horizontal $v_{\rm s}$ component (left), vertical $w_{\rm s}$ component (right). Reference traces for noisy and noiseless data are shown in the top panels.
@@ 4.2 Neural networks
We generated three distinct datasets for the neural network model: a training set consisting of 15,000 samples and a validation set of 3,000 samples. The training and validation datasets were created using fifth-order polynomials and the source wavelet (12) with the DG solver described in Section 3.4, which also details the mesh construction and numerical discretization. The values for the model input $\bm{m}$ were independently drawn from the probability distributions described in Sections 3.2 and 3.3 and used as inputs to the forward model $\mathcal{A}$ (see Eq. (17)). In addition to training and validation datasets, a separate test dataset comprising both synthetic and field measurement data was employed to evaluate the final performance of the neural network. This test dataset is described in detail in Section 5.1.
A fully connected neural network was employed in this study to determine the water volume using seismic data. By training on data with randomized model parameters $\bm{m}$ (see Eq. (14)), the network learns to focus on the relevant seismic signatures of water volume while marginalizing the influence of less significant factors. We down-sample the synthetic data to a sampling frequency of 4 kHz, then generate five copies of the clean training and validation datasets, each corrupted with Gaussian noise according to (15). This augmentation increases the size of both the training and validation datasets by a factor of five, which supports better generalization and robustness to measurement noise. The noise corrupted time domain data are transformed to the frequency domain according to model (17) and denoted as $\hat{X}_{\textrm{noised}}$ in the following. For each sample, we construct four input matrices of size $N_{r}\times N_{f}$, containing the real and imaginary parts of the two velocity components. Here, $N_{r}=56$ is the number of receivers and $N_{f}=21$ is the number of frequencies in the usable-energy band of the data (31.4–88.6 Hz with $\Delta f\approx 2.86$ Hz), see Fig. 5. These four matrices are stacked and reshaped into a single feature vector. This yields an input size of
$\text{frequency components}\times\text{number of receivers}\times\text{velocity components}\times 2=21\times 56\times 2\times 2=4704$
where the factor of 2 at the end accounts for the separate real and imaginary parts. Because the first layer is fully connected, each neuron receives all 4704 features, enabling the model to jointly learn from all four input channels. The network produces a single scalar output representing the estimated water volume.
The training process is implemented using the TensorFlow [1] and Keras [4] libraries. We employ the Adam optimizer [18] and utilize the RandomSearch algorithm from the Keras Tuner library [27] to optimize the model’s performance by exploring different hyperparameter configurations. The Keras Tuner is allowed to randomly select the activation function, learning rate, number of hidden layers, number of neurons per layer, and $L_{2}$ regularization penalty factor. The activation functions considered are “relu,” “sigmoid,” “tanh,” “selu,” “swish,” and “LeakyReLU,” while the number of hidden layers ranges from one to six. The number of neurons per layer varies between 100 and 5,000. The learning rate is selected from $1\times 10^{-3},1\times 10^{-4},1\times 10^{-5}$}, and the $L_{2}$ regularization penalty is chosen from $1\times 10^{-5},1\times 10^{-6},1\times 10^{-7},1\times 10^{-8}$}. For all networks in this study, the batch size is set to 256.
@@
The best-performing network is selected based on MSE for validation. After testing various hyperparameter combinations, the final network consists of five hidden layers with 2,570 (layer 1), 3,920 (layer 2), 3,360 (layer 3), 2,730 (layer 4), and 3,400 (layer 5) neurons. It employs the LeakyReLU activation function, a learning rate of $1\times 10^{-5}$, and an $L_{2}$ regularization penalty factor of $1\times 10^{-5}$. The output layer uses a linear activation function, and early stopping is enabled during both hyperparameter tuning and training to enhance generalization and prevent overfitting.
The training process uses large, randomized datasets that broadly sample the parameter space, with noise added to help improve robustness. Hyperparameters are tuned via a random search, and early stopping is applied to reduce overfitting and support generalization to new data. Figure 6 illustrates the approach for estimating water volume from seismic data. The figure shows the sequential data processing steps, the neural network architecture with five hidden layers, and the training and validation loss curves that highlight the model’s learning and generalization over time.
Figure 6: Graphical illustration of the process for estimating water volume from seismic data. The left side depicts the pre-processing phase, including the deconvolution operation used to normalize the source wavelet effects. The right side presents the neural network-based estimation phase, showing the detailed architecture of the best-performing model. The network consists of five hidden layers with numbers of neurons $(m_{1},m_{2},m_{3},m_{4},m_{5})=(2570,3920,3360,2730,3400)$. Each hidden layer uses the LeakyReLU activation function, while the output layer employs a linear activation. The input size to the network is $n=4704$, corresponding to 2121 frequency components, 22 velocity components, 5656 receivers, and $2$ for the real and imaginary parts. The network produces a single scalar output estimating the water volume. The network was trained with a learning rate of $1\times 10^{-5}$ and an $L_{2}$ regularization penalty factor of $1\times 10^{-5}$. Early stopping was applied during both hyperparameter tuning and training to prevent overfitting and improve generalization. The top-right inset displays the loss curves over epochs for training and validation datasets, illustrating the model’s convergence behavior.
@@ 5.1 Predictions of water volume
We test the applicability of the trained neural network model via two different test datasets. These datasets are defined as:
@@
  • Field Measurements: Field data were acquired at seven distinct water-table levels, ranging from a depth of –31.3 cm to –88.7 cm. At each level, three repeated measurements were recorded for each of three different drop heights, resulting in a dataset comprising $3\times 3\times 7=63$ observations. Using our knowledge of the actual water-table level and the geometry of the sand pool, we can calculate the volume of the water-saturated zone. Additionally, based on a previous study [28], we assume the nominal porosity of the material to be 35 per cent;
@@
  • Synthetic Dataset: This test dataset contains a total of 3,000 samples, distinct from the training (15,000 samples) and validation (3,000 samples) datasets. The same prior distributions were used for material parameter sampling as in the training and validation datasets. Noise parameters $A$ and $B$ in (15) were fixed to 1% each, representing moderate noise levels. The DG simulations used fourth-order basis functions and the Ricker wavelet (see Eq. (13)) as the source function. As noted earlier, a modified mesh density and polynomial basis order were used to introduce controlled numerical variation from the training and validation datasets. Additionally, the Ricker wavelet is employed here, differing from the first derivative of a Gaussian wavelet used in the training phase, to test the neural network’s ability to generalize to unseen source signatures.
@@
The Keras Tuner-optimized network architecture undergoes ten training runs, and the final water volume per sample result is determined as the average of these ten estimates. Figure 7 presents a comparison between the estimated (average) water volumes and their corresponding true values. The results corresponding to the 5 cm, 10 cm, and 15 cm drop heights are shown as three distinct groups, each containing the three repeated measurements $(\rm{dh}_{5}^{1,2,3})$, $(\rm{dh}_{10}^{1,2,3})$, and $(\rm{dh}_{15}^{1,2,3})$, respectively. These drop heights are colour coded in Figure 7, with red corresponding to 5 cm, green to 10 cm, and blue to 15 cm. To get a crude approximation for the uncertainty, we assumed that the porosity value in the field measurements database is uncertain in a sense that we assumed $\phi_{\rm{true}}\in[0.95,1.05]\times 35$ per cent, that can be used to compute the error bars shown for each estimate with real data. The figure shows also the estimates for the synthetic dataset. These results demonstrate the potential of using proposed neural network-based approach to recover the water volume.
@@
Figure 7: Left: Estimated volumes of water as a function of the true value. Red denotes repeated measurements from drop height 5 cm ($\rm{dh}_{5}^{1,2,3}$), green from drop height 10 cm ($\rm{dh}_{10}^{1,2,3}$), and blue from drop height 15 cm ($\rm{dh}_{15}^{1,2,3}$). Right: Histogram of relative prediction errors, with the colour coding of the asterisks corresponding to the same drop heights as shown on the left.
These correlations are key to verifying the accuracy of our neural network model and demonstrate its potential for practical application. Quantitative comparison between seismic data-derived estimates and ground truth water volumes confirms the model’s reliability under controlled conditions. However, the occurrence of a significantly biased estimate for one field sample highlights the challenge posed by out-of-training-distribution data, likely caused by extreme or unexpected measurement noise, potentially hidden in the source-to-receiver offsets. This suggests the importance of sensitivity analysis and uncertainty quantification in further refining the method.
@@
The results at water-table level -36.2 cm reveal a clear outlier in Fig. 7. To analyze the differences between repeated measurements, we utilize the RMSE. Let $d_{i}^{k}$ represent the normalized seismic data (see model (17)) for the $k$’th repeated measurement at drop heights $i\in\{5,10,15\}$ in the frequency domain, with the real and imaginary components stacked. The root mean square error (RMSE) error can now be expressed as follows:
@@
Table 2: Root mean square errors (19) for different data combinations for measurements taken at a water-table level of -36.2 cm.
@@ 5.2 SHAP analysis
The evaluation metrics indicate that the full receiver array provides the highest estimation accuracy, while SHAP value-based selection offers a practical compromise with somewhat reduced accuracy. Random receiver selection significantly degrades performance, underscoring the importance of receiver placement and data quality in seismic monitoring. These findings emphasize the value of explainability methods such as SHAP in guiding sensor array design and improving robustness of neural network-based estimations in real-world conditions.
6 Conclusions
In this study, we investigated the estimation of water volume in a reservoir using both field and synthetic seismic data. The reservoir’s controllable water-table level and known physical characteristics enabled direct validation of our neural network-based estimation approach. Synthetic training data were generated using a coupled poroviscoelastic–viscoelastic wave propagation model solved with a high-order discontinuous Galerkin method on a GPU cluster. The forward simulations used domain-wise homogeneous material properties and allowed for varying water-table levels within the geometry.
A fully connected neural network was trained to estimate water volume from frequency-domain seismic data normalized by a deconvolution method that accounts for unknown source wavelets. Training data were corrupted with noise levels matching field data conditions, ensuring robustness against realistic measurement noise.
The neural network results indicate that, with a full receiver array, water volume can potentially be estimated with good accuracy, approaching the known reservoir volumes. Analysis using SHAP values suggests that, in the present controlled setup, selecting the top ten most influential receivers can maintain performance comparable to the full array, while random receiver selection tends to degrade accuracy, particularly for field data. These findings highlight that receiver contributions influence estimation accuracy in this experiment. However, they are specific to the present setup and should not be assumed to generalize to other geological settings or problems.
Overall, this work presents a promising framework for seismic-based water volume estimation using neural networks. The findings emphasize the value of carefully chosen receiver configurations and suggest that models trained on synthetic data may generalize to controlled real-world conditions. While this study focuses on a simplified and well-constrained setup, the results point to the potential applicability of this approach in more complex environments in future work.